OpenStudy (automaticloveletter):

Which of the following graphs best represents the solution to the pair of equations below? y = −x + 6 y = 2x − 3

OpenStudy (anonymous):

What graphs are there?

OpenStudy (anonymous):

@automaticloveletter

OpenStudy (automaticloveletter):

http://prntscr.com/7lnqmx

OpenStudy (automaticloveletter):

http://prntscr.com/7lnqsh @Tide_Poole

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

http://i.imgur.com/V8Y8M8Q.png

OpenStudy (anonymous):

It's this one

OpenStudy (anonymous):

@automaticloveletter

OpenStudy (automaticloveletter):

which one?

OpenStudy (automaticloveletter):

@Tide_Poole

OpenStudy (anonymous):

Well x=3 so it is going to be the graph where it intersects at (3,3)

OpenStudy (anonymous):

The very first one

OpenStudy (anonymous):

@automaticloveletter

OpenStudy (automaticloveletter):

Oh, thank you. :)

OpenStudy (anonymous):

No problem if you need any help just tag me.

OpenStudy (automaticloveletter):

Sorry I gave the medal to them because they gave the explanation

OpenStudy (automaticloveletter):

Can you help here?

OpenStudy (anonymous):

Sure

OpenStudy (automaticloveletter):

A pair of equations is shown below. x + y = 5 y = one halfx + 2 If the two equations are graphed, at what point do the lines representing the two equations intersect? (2, 5) (5, 2) (2, 3) (3, 2)

OpenStudy (anonymous):

I want you to know something you can figure out the answer in GeoGebra.org it's the easiest way so make two separate lines (x+y=5) and y= (1/2)x+2. So the answer is (2,3). It only takes like 30 second to do it and is really easy.

OpenStudy (automaticloveletter):

Thanks for the help and the website! @alexistheking777